OpenStudy (anonymous):

Evaluate. antilog (-0.0522)?

OpenStudy (anonymous):

\[\frac{ 1 }{ 10^{0.0522}} = 0.8867476\]

OpenStudy (jhannybean):

Can the fraction part not be written as a negative exponent?

OpenStudy (jhannybean):

\[\log (x^{-1}) = \log\left(\frac{1}{x}\right)\]

OpenStudy (anonymous):

^ yeah that makes sense